# FlashDMoE: Fast Distributed MoE in a Single Kernel

### v2

@@ 3.1 Task Abstraction for Computation
Computational operators. The FFN operator is a standard position-wise feed-forward network widely used in Transformer architectures [5], composed of two linear transformations separated by a nonlinear activation $\phi$ (e.g., GELU or ReLU):
@@
Here, $W_{1}$ and $W_{2}$ represent learnable weight matrices, and $b_{1}$ and $b_{2}$ are biases. The expert-combine operation, used in architectures like GShard [33] and DeepSeek [1], merges outputs from multiple experts by computing a weighted combination based on their affinity scores:
@@
The operator $\star_{t}$ (instantiated from $\star$) may behave differently depending on the task metadata $\mathcal{M}$, and the result of $A\star_{t}B$ is accumulated into $D$. We provide an example of task metadata in §D.
@@ 3.2 Symmetric Tensor Layout for Inter-GPU Communication
(a) Symmetric Tensor Layout across 2 Expert-parallel Processes.
@@
Within a single GPU device, the actors in FlashDMoE communicate through the GPU’s memory subsystem (see Figure 8). Specifically, the Scheduler and Subscriber actors exchange data via fast shared memory, while other actor pairs communicate through global memory. For communication across multiple devices, FlashDMoE uses device-initiated communication, leveraging the one-sided PGAS (Partitioned Global Address Space) programming model [34]. However, achieving scalable and correct one-sided memory accesses in PGAS without costly synchronization is a known challenge [1, 35]. We address this challenge with a provably correct and scalable solution: a symmetric tensor layout $L$, supporting fully non-blocking memory accesses. We define L as:
@@
where: $P$ is the expert parallel world size, $R$ identifies communication rounds (i.e., two rounds, one for token dispatch and one for combine), $B$ is number of staging buffers, $E$ is the number of local experts, $C$ is the upscaled expert capacity (§3.2.1) and $H$ is the token embedding dimension. Our core insight to enable non-blocking communication is temporal buffering. Specifically, we overprovision memory for the underlying token matrix by at least $2\cdot r$ times, where $r$ is the number of communication rounds in the dependency graph, and the factor of $2$ accounts for separate buffers for incoming and outgoing data within each communication round. For MoE models, we have $42\cdot r=$ 4. This modest increase in memory usage eliminates the need for synchronization during one-sided data transfers. Figure 9(b) illustrates how cells within this symmetric tensor layout are indexed and used for Direct Memory Access (DMA) and Remote DMA (RDMA) operations. As Theorem 3.1 reinforces, this indexing scheme over $L$ is the underlying mechanism that allows for fully non-blocking accesses eliding synchronization because all accesses are write conflict-free. See§ C for the proof.
@@
$C\cdot E\cdot W=C\cdot E^{\prime}=S^{\prime},\quad\text{where}\quad S^{\prime}\geq S\text{ and }E^{\prime}\geq E_{W}$
In the typical case of uniform expert distribution (illustrated in Figure 9(a)), we have $S^{\prime}=S$ and $E^{\prime}=E_{W}$, where $E_{W}$ is the total number of experts in the model. Thus, the size of the token buffer is $Size(T)=S^{\prime}\cdot H$. In Figure 9(a), each cell labeled $E_{i}$ (with $i\in\{0,\ldots,3\}$) is a matrix of size $(C,H)$. Extending prior work [33, 11], we introduce additional temporal dimensions $R$ (communication rounds) and $B$ (staging buffers). Each communication round has two fixed staging slots: one for outgoing tokens and another for incoming tokens. Each slot, indexed by dimension $P$, forms a tensor of shape $(S^{\prime},H)$. Therefore, the tensor size $Size(L)$ is generally at least four times the original token buffer size, becoming exactly four times larger in the case of uniform expert distribution. Empirically, we find:
$Size(L)\approx 4\cdot Size(T)$
@@ 3.2.1 In-place Padding for Payload Efficiency
Due to the dynamic and uneven distribution of tokens in MoE dispatch [36], GPUs commonly receive fewer tokens than their predefined expert capacity. Current MoE frameworks [10] typically pad these buffers with null tokens before computation, unnecessarily increasing communication payloads and degrading performance. In contrast, we propose in-place padding, performing padding directly within the local symmetric tensor buffers and thus eliminating excess network communication.
@@
4 Experiments
We implement (§G) and evaluate FlashDMoE and evaluate across five metrics: Forward Latency (§ 4.2), GPU Utilization (§ 4.3), Overlap Efficiency (§ 4.4), Throughput (§ 4.5), and Expert Scalability (§ 4.6). We run experiments on a server with 8 NVIDIA H100 80G GPUs interconnected via NVLink, 125 GB of RAM, and 20 vCPUs. We used PyTorch 2.6.0, CUDA 12.8, and Ubuntu 22.04. All experiments use MoE transformer models configured with 16 attention heads, an embedding dimension of 2048, and an FFN intermediate size of 2048. We apply Distributed Data Parallelism (DDP) and Expert Parallelism for all experiments. We execute only the forward pass over a single MoE layer and measure the average runtime of 32 passes after 32 warmup passes. We use top-2 routing with a capacity factor of 1.0. We compare FlashDMoE against several state-of-the-art MoE systems: (1) Comet [11], (2) FasterMoE [13], (3) Megatron-CUTLASS [37], and (4) Megatron-TE: Megatron-LM with Transformer Engine [38]. Comet relies on cudaMemcpyPeerAsync [39], while FasterMoE and Megatron-LM use NCCL exclusively for communication. We also evaluate FlashDMoE on a multi-node environment and discuss our findings in §F.
4.1 Desiderata
In our experiments, we observe Comet exhibiting anomalously bad performance values at 8 GPUs, so we exclude their results from evaluations at 8 GPUs and only include for results at $\leq$ 4 GPUs. Note we evaluate FlashDMoE using FP32 precision whereas all baselines use FP16. We do so because (1) no baseline supports FP32 and (2) time constraints prevent us from tuning our system to peak performance at FP16. Most importantly, this precision discrepancy disadvantages FlashDMoE by doubling the communication and computation precision, making our results a conservative lower bound. Yet, as we show in the succeeding sections, FlashDMoE outperforms all baselines.
@@ 4.2 Forward Latency
We first measure the forward latency of FlashDMoE across different sequence lengths on both 4 and 8 GPU setups (Figure 10). FlashDMoE consistently outperforms all baselines, with especially notable improvements at longer sequence lengths. On 4 GPUs, it achieves up to 4.6x speedup over Megatron-TE at 16K tokens, and 2.6x over FasterMoE. The gains are even more pronounced at 8 GPUs where FlashDMoE maintains low latency, exhibiting up to 6.4x speedup over baselines that degrade steeply due to increasing communication costs as token buffers increase proportionally. These results highlight FlashDMoE’s ability to scale token throughput without suffering from the communication penalties that plague other implementations.
@@ 4.3 GPU Utilization
Figure 11: Comparison of SM utilization, defined as the ratio of cycles in which SMs have at least one warp in flight to the total number of cycles [40]. Values represent the average SM utilization over 100 iterations. All experiments use T = 8K and E = 64 on two A100s
To quantify GPU efficiency, we measure Streaming Multiprocessor (SM) utilization during the forward pass (Figure 11). FlashDMoE achieves 93.17% average SM utilization, over 9x higher than FasterMoE (9.67%), 6.8x higher than DeepEP+Megatron-LM (13.55%) 4x higher than Megatron-TE (59.11%), and 2.2x higher than Comet (42.31%). This improvement stems from our fully fused kernel architecture and fine-grained pipelining of compute and communication tasks. By eliminating idle gaps due to kernel launches and enabling in-kernel task scheduling, FlashDMoE ensures SMs remain busy with productive work throughout execution.
@@ 4.4 Overlap Efficiency
(b) Weak scaling efficiency.
Figure 12: Forward Latency as the Number of Tokens per GPU increases. We define Overlap Efficiency $O_{e}$ to be $O_{e}=T(2)/T(N_{G})$, where $T(N_{G})$ is the latency at $N_{G}$ GPUs and $T(2)$ is the latency at 2 GPUs.
We evaluate the extent to which FlashDMoE overlaps communication and computation by measuring weak scaling efficiency as the number of GPUs increases (Figure 12(b)). We note that most baselines fail to execute at a single GPU, hence why we use 2 GPUs as the reference point. We observe that Megatron-CUTLASS and Megatron-TE degrade significantly, with overlap efficiency dropping below 0.5 at $\geq 4$ GPUs. FlashDMoE gives up to 3.88x and 4x higher efficiency at 4 and 8 GPUs, respectively. Figure 12(a) further illuminates this efficiency, as FlashDMoE shows stable forward latency growth, whereas baselines Megatron-CUTLASS and Megatron-TE experience approximately linear latency amplification while FasterMoE exhibits sublinear scaling. We attribute this suboptimal performance to straggler effects and exposed communication. In contrast, FlashDMoE demonstrates uniform latency as expected since the workload per GPU is fixed in this weak scaling experiment. These results further corroborate that FlashDMoE’s actor-based design and asynchronous data movement achieve near-ideal overlap, even at scale.
4.5 Throughput
Figure 13: Throughput as the amount of GPUs increases. We compute throughput as $\frac{T*N_{G}}{latency}$, where $N_{G}$ is the number of GPUs.
Throughput, measured in tokens per second (MTokens/s), reflects end-to-end system efficiency. As shown in Figure 13, FlashDMoE scales linearly with GPU count, reaching 17.7 MTokens/s at 8 GPUs. This is over 5.7x higher than FasterMoE and 4.9x higher than Megatron-TE and Megatron-CUTLASS. Notably, these results are achieved despite FlashDMoE operating entirely in FP32, while baselines use FP16. This indicates that FlashDMoE’s design eliminates throughput bottlenecks not by exploiting lower precision, but by maximizing hardware utilization and eliminating host-driven inefficiencies.
@@ 4.6 Expert Scalability
We analyze how FlashDMoE scales with increasing number of experts at fixed sequence length (T = 16K). Note that for the discussed plots, the number of experts on the x-axis is the total number across all GPUs. Each GPU gets 1/8th of this value. As seen in Figure14, FlashDMoE maintains low, uniform latency, as desired, even as the number of experts grows from 8 to 128. In contrast, baselines exhibit superlinear latency increases due to increased kernel launch overheads. FlashDMoE outperforms these baselines by up to 4X at 4 H100s and 6.6X at 8 H100s, both at 128 experts. FlashDMoE ’s payload-efficient communication and scheduler-driven in-kernel dispatching allow it to sustain expert parallelism without incurring the communication and orchestration penalties seen in other systems. These results reinforce FlashDMoE’s scalability for ultra-sparse MoE configurations.
4.7 Memory Overhead
We measure the GPU memory required for the symmetric tensor $L$ and runtime bookkeeping state of FlashDMoE. Memory overhead depends primarily on the tile size, expert capacity (EC), and the number of experts ($E$). Table 3 summarizes memory overhead under various configurations, confirming that FlashDMoE maintains a modest and predictable memory footprint.
Table 3: Memory overhead of FlashDMoE (tile size $bM=128$), $Size(T)=\text{Tokens}*4KB$).
Tokens Experts EC max(bM, EC) Bookkeeping (MB) $Size(L)$ (MB) Total (MB)
4K 16 256 256 64.57 64.00 128.57
4K 32 128 128 64.55 64.00 128.55
4K 64 64 128 128.90 128.01 256.91
4K 128 32 128 257.96 256.02 513.98
8K 16 512 512 128.95 128.01 256.95
8K 32 256 256 128.90 128.01 256.91
8K 64 128 128 128.90 128.01 256.91
8K 128 64 128 258.15 256.02 514.17
16K 16 1024 1024 257.89 256.02 513.90
16K 32 512 512 257.79 256.02 513.81
16K 64 256 256 257.80 256.02 513.81
16K 128 128 128 258.53 256.02 514.54
@@ 5 Limitations and Future Work
Despite the performance gains and architectural innovations of FlashDMoE, there are several limitations worth acknowledging—both practical and conceptual—that open the door to future research.
  • Programming complexity. Developing fully fused, persistent kernels is a non-trivial engineering task. While FlashDMoE proves the feasibility and benefit of such kernels, their construction demands deep expertise in GPU architectures, synchronization and distributed protocols, and memory hierarchies. This high barrier to entry limits adoption. Future work may consider compiler-level abstractions or DSLs to democratize this technique.
  • FP16 support and shared memory access patterns. Although modern GPUs natively support half-precision computation, adapting FlashDMoE to FP16 is non-trivial for the Processor’s computational operators. Specifically, our manually tuned swizzle shared memory layouts are not the most efficient template parameters for CUTLASS’ Collective Mainloop operator which we use to implement our in-device GEMMs. This suboptimal configuration degrades memory throughput as shown in §H. Overcoming this for Ampere GPUs and below would require careful investigation of optimal layouts, but for Hopper GPUs and above, we anticipate using the builder interface that CUTLASS provides in our future improvements.
  • Lack of backward pass and training support. While this work focuses on inference, enabling training requires fusing backward computation and gradient communication into the kernel. Supporting this entails non-trivial changes to both memory bookkeeping and task descriptor definitions. Nevertheless, it remains an exciting direction for extending this system to fully support end-to-end training.
@@ 6 Conclusion
This work introduces FlashDMoE, the first system to fuse the entire Mixture-of-Experts (MoE) operator into a single, persistent GPU kernel. We show that prevailing MoE implementations suffer from two critical inefficiencies: (1) CPU-managed synchronous communication that leads to underutilized interconnects and (2) fragmented execution via multiple GPU kernels, introducing overhead and synchronization delays.
In contrast, FlashDMoE embraces a model of GPU autonomy by embedding computation, communication, and scheduling within a unified kernel. It leverages actor-style concurrency, warp specialization, and asynchronous (R)DMA to achieve fine-grained communication–computation overlap.
Our evaluation demonstrates up to 6$\times$ speedup over state-of-the-art systems, up to 9$\times$ improved GPU utilization, and 5.7$\times$ increased throughput for Distributed MoE. FlashDMoE challenges the dominant execution paradigms in distributed deep learning and presents a compelling template for building future GPU-native systems.
While several limitations remain, programming complexity and lack of FP16 support, this work lays the groundwork for a new era of in-kernel distributed computation. Future systems may build upon this foundation to enable kernel fusion for entire training pipelines, ushering in a design shift from CPU orchestration to fully autonomous GPU execution.

### v3

@@ 3.1 Task Abstraction for Computation
Computational operators. The FFN operator is a standard position-wise feed-forward network widely used in Transformer architectures [NIPS2017_3f5ee243], composed of two linear transformations separated by a nonlinear activation $\phi$ (e.g., GELU or ReLU):
@@
Here, $W_{1}$ and $W_{2}$ represent learnable weight matrices, and $b_{1}$ and $b_{2}$ are biases. The expert-combine operation, used in architectures like GShard [DBLP:conf/iclr/LepikhinLXCFHKS21] and DeepSeek [deepep], merges outputs from multiple experts by computing a weighted combination based on their affinity scores:
@@
The operator $\star_{t}$ (instantiated from $\star$) may behave differently depending on the task metadata $\mathcal{M}$, and the result of $A\star_{t}B$ is accumulated into $D$. We provide an example of task metadata in §E.
@@ 3.2 Symmetric Tensor Layout for Inter-GPU Communication
(a) Layout across 2 Expert-parallel Processes.
@@
Figure 7: Symmetric Tensor Layout
Within a single GPU device, the actors in FlashMoE communicate through the GPU’s memory subsystem. Specifically, the Scheduler and Subscriber actors exchange data via fast shared memory, while other actor pairs communicate through global memory. For communication across multiple devices, FlashMoE uses device-initiated communication, leveraging the one-sided PGAS (Partitioned Global Address Space) programming model [10.1145/1278177.1278183]. However, achieving scalable and correct one-sided memory accesses in PGAS without costly synchronization is a known challenge [deepep, triton-dist]. We address this challenge with a provably correct and scalable solution: a symmetric tensor layout $L$, supporting fully non-blocking memory accesses. We define L as:
@@
where: $P$ is the expert parallel world size, $R$ identifies communication rounds (i.e., two rounds, one for token dispatch and one for combine), $B$ is number of staging buffers, $E$ is the number of local experts, $C$ is the upscaled expert capacity (§3.2.1) and $H$ is the token embedding dimension. Our core insight to enable non-blocking communication is temporal buffering. Specifically, we overprovision memory for the underlying token matrix by at least $2\cdot r$ times, where $r$ is the number of communication rounds in the dependency graph, and the factor of $2$ accounts for separate buffers for incoming and outgoing data within each communication round. For MoE models, we have $2\cdot r=4$. This modest increase in memory usage eliminates the need for synchronization during one-sided data transfers. Figure 7(b) illustrates how cells within this symmetric tensor layout are indexed and used for Direct Memory Access (DMA) and Remote DMA (RDMA) operations. As Theorem 3.1 reinforces, this indexing scheme over $L$ is the underlying mechanism that allows for fully non-blocking accesses eliding synchronization because all accesses are write conflict-free. See§ C for the proof.
@@
[…]
In the typical case of uniform expert distribution (illustrated in Figure 7(a)), we have $S^{\prime}=S$ and $E^{\prime}=E_{W}$, where $E_{W}$ is the total number of experts in the model. Thus, the size of the token buffer is $Size(T)=S^{\prime}\cdot H$. In Figure 7(a), each cell labeled $E_{i}$ (with $i\in\{0,\ldots,3\}$) is a matrix of size $(C,H)$. Extending prior work [DBLP:conf/iclr/LepikhinLXCFHKS21, comet], we introduce additional temporal dimensions $R$ (communication rounds) and $B$ (staging buffers). Each communication round has two fixed staging slots: one for outgoing tokens and another for incoming tokens. Each slot, indexed by dimension $P$, forms a tensor of shape $(S^{\prime},H)$. Therefore, the tensor size $Size(L)$ is generally at least four times the original token buffer size, becoming exactly four times larger in the case of uniform expert distribution. Empirically, we find $Size(L)\approx 4\cdot Size(T)$, contributing memory overhead $\leq 2$% of memory capacity for inference of popular models. We present a thorough breakdown in §D.
@@ 3.2.1 In-place Padding for Payload Efficiency
Due to the dynamic and uneven distribution of tokens in MoE dispatch [bmamba], GPUs commonly receive fewer tokens than their predefined expert capacity. Current MoE frameworks [pmlr-v162-rajbhandari22a] typically pad these buffers with null tokens before computation, unnecessarily increasing communication payloads and degrading performance. In contrast, we propose in-place padding, performing padding directly within the local symmetric tensor buffers and thus eliminating excess network communication.
@@
4 Evaluation
We implement (§G) and evaluate FlashMoE across five metrics: Forward Latency (§ 4.1), GPU Utilization (§ 4.2), Overlap Efficiency (§ 4.4), Throughput (§ 4.3), and Expert Scalability (§ 4.5). We run experiments on a server with 8 NVIDIA H100 80G GPUs interconnected via NVLink, 125 GB of RAM, and 20 vCPUs. We used PyTorch 2.6.0, CUDA 12.8, and Ubuntu 22.04. All experiments use MoE transformer models configured with 16 attention heads, an embedding dimension of 2048, and an FFN intermediate size of 2048. We apply Distributed Data Parallelism (DDP) and Expert Parallelism for all experiments. We execute only the forward pass over a single MoE layer and measure the average runtime of 32 passes after 32 warmup passes. We use top-2 routing with a capacity factor of 1.0. We compare FlashMoE against several state-of-the-art MoE systems: (1) Comet [comet], (2) FasterMoE [fastermoe], (3) Megatron-CUTLASS [megatron], and (4) Megatron-TE: Megatron-LM with Transformer Engine [transformer-engine]. Comet relies on cudaMemcpyPeerAsync [fluxp2p], while FasterMoE and Megatron-LM use NCCL exclusively for communication.
Desiderata. We observe Comet exhibiting anomalously bad performance values at 8 GPUs, so we exclude their results from evaluations at 8 GPUs and only include for results at $\leq$ 4 GPUs. We evaluate FlashMoE using FP32 precision whereas all baselines use FP16. We do so because (1) of incomplete fp16 tuning (§H) and (2) no baseline supports FP32. Note, this precision discrepancy disadvantages FlashMoE by doubling its communication volume and computational workload.
@@ 4.1 Forward Latency
We first measure the forward latency of FlashMoE across different sequence lengths on both 4 and 8 GPU setups (Figure 8). FlashMoE consistently outperforms all baselines, with especially notable improvements at longer sequence lengths. On 4 GPUs, it achieves up to 4.6x speedup over Megatron-TE at 16K tokens, and 2.6x over FasterMoE. The gains are even more pronounced at 8 GPUs where FlashMoE maintains low latency, exhibiting up to 6.4x speedup over baselines that degrade steeply due to increasing communication costs as token buffers increase proportionally.
@@ 4.2 GPU Utilization
Figure 9: SM utilization, defined as the ratio of cycles in which SMs have at least one warp in flight to the total number of cycles [nsight-metrics]. Values represent the average SM utilization over 100 iterations.
To quantify GPU efficiency, we measure Streaming Multiprocessor (SM) utilization during the forward pass (Figure 9). FlashMoE achieves 93.17% average SM utilization, over 9x higher than FasterMoE (9.67%), 6.8x higher than DeepEP+Megatron-LM (13.55%) 4x higher than Megatron-TE (59.11%), and 2.2x higher than Comet (42.31%). This improvement stems from our fully fused kernel architecture and fine-grained pipelining of compute and communication tasks. By eliminating idle gaps due to kernel launches and enabling in-kernel task scheduling, FlashMoE ensures SMs remain busy with productive work throughout execution.
4.3 Throughput
Figure 10: Throughput when scaling the number of GPUs, computed as $\frac{T\times N_{G}}{\text{latency}}$.
As shown in Figure 10, FlashMoE scales linearly with GPU count, reaching 17.7 MTokens/s at 8 GPUs. This is over 5.7x higher than FasterMoE and 4.9x higher than Megatron-TE and Megatron-CUTLASS. Notably, these results are achieved despite FlashMoE operating entirely in FP32, while baselines use FP16. This indicates that FlashMoE ’s design eliminates throughput bottlenecks not by exploiting lower precision, but by maximizing hardware utilization and eliminating host-driven inefficiencies.
@@ 4.4 Overlap Efficiency
(b) Weak scaling efficiency
Figure 11: Weak scaling efficiency. We define Overlap Efficiency $O_{e}$ to be $O_{e}=T(2)/T(N_{G})$, where $T(N_{G})$ is the latency at $N_{G}$ GPUs and $T(2)$ is the latency at 2 GPUs.
We evaluate the extent to which FlashMoE overlaps communication and computation by measuring weak scaling efficiency as the number of GPUs increases (Figure 11(b)). We note that most baselines fail to execute at a single GPU, hence why we use 2 GPUs as the reference point. We observe that Megatron-CUTLASS and Megatron-TE degrade significantly, with overlap efficiency dropping below 50% at $\geq 4$ GPUs. FlashMoE gives up to 3.88x and 4x higher efficiency at 4 and 8 GPUs, respectively. Figure 11(a) further illuminates this efficiency, as FlashMoE shows stable forward latency growth. These results corroborate that FlashMoE’s actor-based design and asynchronous data movement achieve near-ideal overlap.
@@ 4.5 Expert Scalability
We analyze how FlashMoE scales with increasing number of experts at fixed sequence length (T = 16K). Note that for the discussed plots, the number of experts on the x-axis is the total number across all GPUs. Each GPU gets 1/8th of this value. As seen in Figure 12, FlashMoE maintains low, uniform latency, as desired, even as the number of experts grows from 8 to 128. In contrast, baselines exhibit superlinear latency increases due to increased kernel launch overheads. FlashMoE outperforms these baselines by up to 4X at 4 H100s and 6.6X at 8 H100s, both at 128 experts. FlashMoE ’s payload-efficient communication and scheduler-driven in-kernel dispatching allow it to sustain expert parallelism without incurring the communication and orchestration penalties seen in other systems. These results reinforce FlashMoE ’s scalability for ultra-sparse MoE configurations.
@@ 5 Limitations and Future Work
Engineering complexity. Fully fused, persistent kernels demand deep GPU + distributed-systems expertise; future work may investigate compiler/DSL abstractions to lower this barrier.
FP16 inefficiency. Our FP16 path is suboptimal (§H) due to insufficient tuning. We anticipate addressing this gap with autotuned GEMM operators like cuBLASDx [cudx] or CUTLASS builders.
Training support. This work targets inference; enabling training will require fusing backward computation and gradient communication with new bookkeeping and task descriptors.
@@ 6 Conclusion
We introduce FlashMoE, the first work to fuse the entire Distributed MoE operator into a single persistent GPU kernel that unifies computation, communication, and scheduling via actor-style concurrency, warp specialization, and async (R)DMA. We address two dominant bottlenecks in prior systems—CPU-managed synchronous communication and fragmented multi-kernel execution. Empirically, FlashMoE achieves up to speedup, higher GPU utilization, and 5.7× throughput for distributed MoE. Looking ahead, we see a shift from CPU orchestration to fully autonomous, GPU-native pipelines—extending this fusion approach to training and beyond.